Find the area under the standard normal to the right of z = -1.50 and to the left of z = 2.50 (Round answer to the nearest ten-thousandths.)
Answer #1

Area to the right of z = - 1.50, = 1 - 0.0668 = 0.9332

Area to the left of z = 2.50, = 0.9938

Area between them = 0.9938 - 0.0668

= 0.9270